OpenStudy (anonymous):

Solve 4r2 – 9 = 91. A. {± 3} B. {± 4} C. {± 5} D. {±6}

Vocaloid (vocaloid):

add 9 to both sides divide both sides by 4 take the square root of each side

OpenStudy (anonymous):

4r^2=100

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

alright, so you added 9 to each side what's the next step?

OpenStudy (anonymous):

r^2=25

Vocaloid (vocaloid):

good, what's the next step?

OpenStudy (anonymous):

r^2=5

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

not quite, check the left side

OpenStudy (anonymous):

im not sure what to do to the left side!

Vocaloid (vocaloid):

hint: the square root of a^2 = a so what is the square root of r^2?

OpenStudy (anonymous):

r=5

Vocaloid (vocaloid):

you're very close! when we take the square root of something, the answer can be either positive or negative, which is why